OpenStudy (anonymous):

find the lcd 7/t, 6/t-1

OpenStudy (anonymous):

t(t-1)... they have nothing in common that would be like if t = 5 then (t-1) = 4 and the common denominator would be 5 * 4 = 20

OpenStudy (anonymous):

ok so becasue they have nothing in common i keep the t and (t-1) ?

OpenStudy (anonymous):

Yes it is "t" times "t-1" = t(t-1)

OpenStudy (anonymous):

awesomeness :D Gracias (thank you)